OpenStudy (anonymous):

Solve y=-x^2+12x-22 using the quadratic formula. I have an answer, but I want to double check because this is not my strongest chapter in the course.

OpenStudy (amistre64):

well, the quadrtic formula is used for 0 = ax^2 + bx + c y=-x^2+12x-22 0= - x^2 + 12x + (-22-y) \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] \[x=\frac{-12\pm \sqrt{144-4(-1)(-22-y)}}{2(-1)}\]

OpenStudy (anonymous):

\[y=6 \pm \sqrt{14}\] This is what I got. Is it right? @amistre64